# An unexpected rational blowdown

### v1

@@ Abstract.
The rational blowdown operation in 4-manifold topology replaces a neighborhood of a configuration of spheres by a rational homology ball. Such configurations typically arise from resolutions of surface singularities that admit rational homology disk smoothings. Conjecturally, all such singularities must be weighted homogeneous and belong to certain specific families: Stipsicz–Szabó–Wahl constructed QHD smoothings for these families and used Donaldson’s theorem to obtain very restrictive necessary conditions on the resolution graphs for singularities with this property. In particular, these results, as well as subsequent work of Bhupal–Stipsicz, show that for certain resolution graphs, the canonical contact structure on the link of the singularity cannot admit a QHD symplectic filling. By contrast, we exhibit Stein rational homology disk fillings for the contact links of an infinite family of rational singularities that are not weighted homogeneous, producing a new symplectic rational blowdown. Inspiration for our construction comes from de Jong–van Straten’s description of Milnor fibers of sandwiched singularities; we use the symplectic analog of de Jong–van Straten theory developed by the second and third authors. The unexpected Stein fillings are built using spinal open books and nearly Lefschetz fibrations.
@@ 1. Introduction
From another perspective, understanding the topology of Milnor fibers is an interesting question in singularity theory. By [Wah81], a surface singularity must be rational if it admits a smoothing that is a rational homology disk. However, few rational singularities have such smoothings. All known examples fall into several specific families of weighted homogeneous singularities; the corresponding resolution graphs have one node, of valency three or four, [SSW08]. It has been conjectured that the known examples give a complete list of surface singularities with $\mathbb{Q}$HD smoothings.
In [SSW08], Stipsicz–Szabó–Wahl used Donaldson’s theorem to attack the question by examining possible embeddings of the homology lattice of the plumbing into the diagonal lattice of the same rank. This gives very restrictive necessary conditions for the dual resolution graphs whose link can admit a symplectic $\mathbb{Q}$HD filling. Bhupal–Stipsicz ([BS11]) settled additional cases by studying symplectic fillings of the links of the corresponding singularities with their canonical contact structures: they showed that the only weighted homogeneous singularities whose links admit $\mathbb{Q}$HD fillings are those with known $\mathbb{Q}$HD smoothings. (This classification is reproven and extended with methods closer to the ones used in this paper in [Bek25] when combined with the results of [PS25].)
@@
Figure 1. The family $G_{k,n}$ for $k\geq-1,n\geq 0$, where $G_{k,n}$ has $k+n+7$ vertices, with $k$ vertices between the two nodes. The graph $G_{-1,1}$ has a single node of valency 44, with 6 vertices total.
For $k\geq 0$, the graphs of Figure 1 all have two nodes and belong to the family $\mathcal{C}$ of [SSW08], producing a new symplectic rational blowdown. It is interesting to note that for [^1] This notation means contracting the edge between the two nodes, and giving them framing $a+b+2$ if the two nodes were framed $a,b$ respectively. k=1k=-1, a rational homology disk smoothing actually exists for a singularity in the corresponding topological type by [Fow13, Theorem 5(a)], when the cross ratio of the intersection points of the node is $9$. Our construction produces a $\mathbb{Q}$HD Stein filling in this case as well, but we do not know whether our filling is diffeomorphic to the Milnor fiber. With this in mind, our construction is an extension of the $\mathcal{C}^{4}$ family of [BS11].
@@
These graphs all correspond to sandwiched singularities (in combinatorial terms, this means that the graph can be augmented, by adding (1)(-1) vertices, to a plumbing graph that can be blown down to a smooth point). Deformation theory for this class of singularities can be understood thanks to a very attractive approach of de Jong–van Straten [dJvS98]: the surface singularity $(X,0)$ is encoded via the germ of a singular plane curve $(C,0)$, and all smoothings of $(X,0)$ correspond to certain picture deformations of $(C,0)$, decorated with marked points. Milnor fibers can be reconstructed directly from picture deformations. In particular, a Milnor fiber of a smoothing is a rational homology disk if and only if the number of irreducible components of $(C,0)$ matches the number of marked points of the corresponding picture deformation. The combinatorial features of the arrangement of curves and marked points in a picture deformation depend on the original dual resolution graph. In [Bek25], the first author was able to rule out rational homology disk smoothings for certain families of graphs through combinatorial analysis of possible curve and point arrangements. The family of Figure 1 was discovered in the process: combinatorial obstructions vanish, even though rational homology smoothings do not exist in this case by [Wah11, Theorem 8.6], since the graphs of Figure 1 are taut of type $(L_{2})-(J_{1})-(R_{1})$ in the sense of Laufer [Lau73, Section 2.2]. For symplectic fillings of the contact link of a sandwiched singularity, the second and third author developed an analog of the de Jong–van Straten theory [PS23, PS25], showing that all fillings arise in a similar way from certain immersed disk arrangements in $\mathbb{C}^{2}$ compatible with the germ $(C,0)$ associated to the singularity. The constructions of [PS25] use spinal open books and nearly Lefschetz fibrations; the symplectic input comes from [LHMW18, LHMW20, BH16, MRW24].
@@
From the de Jong–van Straten correspondence and non-existence of the corresponding $\mathbb{Q}$HD smoothings, it follows that the combinatorial arrangement producing our $\mathbb{Q}$HD filling cannot be realized by an analytic deformation of the corresponding curve germ. We do not know how to prove this deformation non-realizability by a direct argument for analytic deformation of plane curves; it would be interesting to investigate this phenomenon further.
@@ 2. Picture deformations, DJVS arrangements, and $4$-manifold constructions
A normal surface singularity $(X,0)$ is sandwiched if there exists an embedding of the tubular neighborhood of the exceptional set of its resolution $\tilde{X}$ into some blow-up of $\mathbb{C}^{2}$. One can consider arbitrary resolutions here; for simplicity, we always work with the minimal good resolution, where the exceptional curves are smooth and intersect at transverse double points only. (Note that the minimal resolution is good for all rational singularities.) A good resolution can be encoded by the dual resolution graph, whose vertices correspond to the irreducible components of the exceptional divisor, and the edges record intersections between the components. Sandwiched singularities can be defined by a combinatorial condition on the dual resolution graph: a graph is sandwiched if it can be augmented by adding new end vertices with self-intersection (1)(-1) in a way that the resulting augmented graph can be blown to a smooth point at 0. (The choice of such augmentation is generally not unique.) The (1)(-1) vertices correspond to a distinguished collection of (1)(-1) curves in a blow-up of $\mathbb{C}^{2}$ with the embedding of $\tilde{X}$, so that the configuration of these (1)(-1) curves together with the original exceptional set can be completely blown down.
In [dJvS98], the sandwiched singularity is then encoded by an associated singular plane curve germ, as follows. For each distinguished (1)(-1) curve, fix a transverse complex disk $\tilde{C_{i}}$ through a generic point. Blow down the configuration of the exceptional curves of $(X,0)$ together with these (1)(-1) curves; the image $C_{i}$ of the disk $\tilde{C_{i}}$ under the blow-down is (a germ of) a curve in $\mathbb{C}^{2}$, possibly singular at the origin. We set $C=\cup C_{i}$ and decorate $C$ with a weight $w$, a collection of integers $w_{i}=w(C_{i})$. For each $C_{i}$, the number $w_{i}$ is the sum of multiplicities of the intersections of the image of $\tilde{C_{i}}$ with the exceptional curves during the blow-down process. (If $C_{i}$ is smooth, then $w_{i}$ is simply the number of blow-downs that the corresponding component goes through.) The original singularity can be reconstructed from the decorated germ $(C,w)$, where $C=\cup C_{i}$ and $w$ is collection of weights $w_{i}$: the corresponding blow-ups recover the embedding of the tubular neighborhood of the exceptional set of the resolution given by $G$ into some blow-up of $\mathbb{C}^{2}$.
@@
Figure 2 shows the augmented graph, the blow-down process, and the resulting decorated germ for the graph $G_{0,1}$ of Figure 1. Each component $C_{i}$ of the germ is modelled on the cusp $y^{2}=x^{3}$ because the previous steps of the blowdown give a familiar embedded resolution of the cuspidal cubic. We also see that components $C_{i}$ and $C_{j}$ have intersection multiplicity $8$ when $i,j=1,\dots,6$, and $C_{0}$ and $C_{i}$ have multiplicity of intersection $7$ for each $i=1,\dots,6$. In general, for the graphs $G_{k,n}$ with $k\geq 0$ we get cusps $C_{1},...,C_{6+k}$ with decoration $7+k$ and pairwise intersection multiplicity $8+k$ for each pair, and instead of $C_{0}$ we have $n$ cusps $C^{\prime}_{1},\dots C^{\prime}_{n}$ with decoration $7+k+n$, with $C^{\prime}_{i}\cdot C^{\prime}_{j}=8+k+n$ for $1\leq i<j\leq n$ and $C^{\prime}_{j}\cdot C_{i}=7+k$ for $i=1,\dots,k+6$, $j=1,\dots,n$.
@@
Strictly speaking, our pictures encode topological rather than analytic data. If we consider the exceptional set of the resolution $\tilde{X}$ as an embedded configuration of complex curves, then, once the augmented graph is fixed, $C$ is a germ of a reducible complex curve in $\mathbb{C}^{2}$ (defined up to analytic equivalence). However, if we only fix the resolution graph and its augmentation, the output of the blowdown process is the topological type of a germ $C$, that is, the link $C\cap S^{3}$ of the plane curve singularity $(C,0)$ obtained by intersecting the curve $C$ with a small sphere centered at the origin. (This is a link in $S^{3}$ in the knot-theoretic sense, defined up to smooth isotopy.) By [CPP04], the contact link $(Y,\xi)$ of the singularity $(X,0)$ depends only on the topological type; to generate symplectic fillings, we will only need to know the isotopy class of $C\cap S^{3}$.
@@
(C-3) the marked points $p_{j}$ are chosen on $C^{t}$, so that each intersection between two irredicible components is marked; additionally, there may be free marked points on each $C_{i}$ away from the intersections,
@@
In [dJvS98], $p$ is the deformation of the corresponding scheme, originally concentrated at 0 and reduced for $t>0$, but we will think of the marked points simply as a set.
The $\delta$-constant property implies that the deformation preserves the branches, and each $C^{t}_{i}$ is an immersed disk in $B$, [Tei76]. With the appropriate choice of coordinates, we can assume that none of the tangent cones of the branches $C_{i}$ of the decorated germ are vertical and, moreover, all vertical tangencies of the projections $C^{t}_{i}\to\mathbb{C}_{x}$ are nondegenerate. This means that outside of self-intersections, each projection is a branched covering with simple branch points only; the degree of the covering is the multiplicity of the corresponding branch $C_{i}$ at 0. We fix a good representative of the deformation in a Milnor ball, thought of as the product $B=D_{x}\times D_{y}$ of two coordinate disks, with corners smoothed. All the constructions will take place inside this ball. We will also assume that $C\cap\partial B$ is contained in $\partial D_{x}\times D_{y}$, and think of the link of $C$ as a braid.
@@
($\Gamma$-2) All intersections $\Gamma_{i}\cap\Gamma_{k}$ and self-intersections of the components $\Gamma_{i}$ are positive transverse multiple points locally modelled on the intersection of complex lines.
($\Gamma$-3) All intersections are marked, and there can be additional free marked points.
@@
[dJvS98] Every Milnor fiber of a deformation of a sandwiched singularity $(X,0)$ can be obtained by the DJVS construction from some picture deformation of a fixed decorated germ $(C,w)$ for $(X,0)$.
@@
[PS25] (1) For every DJVS arrangement $(\Gamma,p)$ compatible with $(X,0)$, the 4-manifold $W_{(\Gamma,p)}$ carries a Stein structure and gives a Stein filling of the contact link $(Y,\xi)$ of $(X,0)$.
@@
Indeed, it is not hard to see that $H_{1}(W_{\Gamma,p})$ is generated by loops around $\tilde{\Gamma}_{i}$’s, $H_{2}(W)$ can be identified with the kernel of the incidence map $\mathcal{I}:\mathbb{Z}\langle{p_{j}}\rangle\to\mathbb{Z}\langle{\Gamma_{i}}\rangle$, and $H_{1}(W)$ is the cokernel of this map. We have $b_{1}=0$ because $W$ consists of 1- and 2-handles only, and $b_{1}(Y)=0$ for the link $Y=\partial W$ since the singularity is rational, c.f. [dJvS98, Theorem 5.2], [PS23, Section 6.1]. Thus, to produce a $\mathbb{Q}$HD filling, the arrangement $(\Gamma,p)$ must have the number of marked points matching the number of components $\Gamma_{i}$; this is rare, since typically an arrangement $\Gamma$ has a much larger number of marked points than components.
To illustrate the discussion, we describe the Milnor fiber of one of the smoothings of the singularities in our family. Consider the germ $C$ of Figure 2 and take its Scott deformation, constructed iteratively as follows. We blow up at the singular point at the origin, perform a small deformation of the strict transform so that its singularities become disjoint from the exceptional curve $E$, and the branches of the strict transforms are transverse to $E$ (equivalently, we can think of this as “shifting” the exceptional curve $E$ off the singularity in the strict transform), and then blow down $E$ to get a deformed curve $C^{\prime}$ in $\mathbb{C}^{2}$. As a result, the curve $C^{\prime}$ has a new transverse multiple point as well as the collection of singularities occurring on the strict transform of $C$ in the blow-up of $\mathbb{C}^{2}$ at $p$. The procedure is repeated until all the singularities at the given stage are transverse multiple points. See [dJvS98, A’C75] for details, including the explanation why this procedure can be actually realized by a 1-parameter deformation. Figure 3 illustrates the Scott deformation of our decorated germ from Figure 2. The marked points on the picture deformation include all the intersections as well as the free points to match the weights on the components of the singular germ $C$. (Decorated germs look similar for all singularities corresponding to graphs $G_{0,n}$, with the number of intersection points increasing with $n$, although the parity of $n$ affects the order of curve “tails” on the right side of the picture.) For an arbitrary decorated germ, the Scott deformation produces the Artin smoothing under the de Jong–van Straten correspondence, with the Milnor fiber that is diffeomorphic to the minimal resolution of the surface singularity [dJvS98].

### v2

@@ Abstract.
The rational blowdown operation in 4-manifold topology replaces a neighborhood of a configuration of spheres by a rational homology ball. Such configurations typically arise from resolutions of surface singularities that admit rational homology disk smoothings. Conjecturally, all such singularities must be weighted homogeneous and belong to certain specific families: Stipsicz–Szabó–Wahl constructed $\mathbb{Q}$HD smoothings for these families and used Donaldson’s theorem to obtain very restrictive necessary conditions on the resolution graphs for singularities with this property. In particular, these results, as well as subsequent work of Bhupal–Stipsicz, show that for certain resolution graphs, the canonical contact structure on the link of the singularity cannot admit a $\mathbb{Q}$HD symplectic filling. By contrast, we exhibit Stein rational homology disk fillings for the contact links of an infinite family of rational singularities that are not weighted homogeneous, producing a new symplectic rational blowdown. Inspiration for our construction comes from de Jong–van Straten’s description of Milnor fibers of sandwiched singularities; we use the symplectic analog of de Jong–van Straten theory developed by the second and third authors. The unexpected Stein fillings are built using spinal open books and nearly Lefschetz fibrations.
@@ 1. Introduction
From another perspective, understanding the topology of Milnor fibers is an interesting question in singularity theory. By [WAH81], a surface singularity must be rational if it admits a smoothing that is a rational homology disk. However, few rational singularities have such smoothings. All known examples fall into several specific families of weighted homogeneous singularities; the corresponding resolution graphs have one node, of valency three or four, [SSW08]. It has been conjectured ([WAH11, Conjecture 8.10]) that the known examples give a complete list of surface singularities with $\mathbb{Q}$HD smoothings.
In [SSW08], Stipsicz–Szabó–Wahl used Donaldson’s theorem to attack the question by examining possible embeddings of the homology lattice of the plumbing into the diagonal lattice of the same rank. This gives very restrictive necessary conditions for the dual resolution graphs whose link can admit a symplectic $\mathbb{Q}$HD filling. Bhupal–Stipsicz ([BS11]) settled additional cases by studying symplectic fillings of the links of the corresponding singularities with their canonical contact structures: they showed that the only weighted homogeneous singularities whose links admit $\mathbb{Q}$HD fillings are those with known $\mathbb{Q}$HD smoothings. (This classification is reproved and extended with methods closer to the ones used in this paper in [BEK25] when combined with the results of [PS25].)
@@
Figure 1. The family $G_{k,n}$ for $k\geq-1,n\geq 0$, where $G_{k,n}$ has $k+n+6$ vertices, with $k$ vertices between the two nodes. The graph $G_{-1,1}$ has a single node of valency 44, with 6 vertices total.
For $k\geq 0$, the graphs of Figure 1 all have two nodes and belong to the family $\mathcal{C}$ of [SSW08], producing a new symplectic rational blowdown. It is interesting to note that for [^1] This notation means contracting the edge between the two nodes, and giving them framing $a+b+2$ if the two nodes were framed $a,b$ respectively. k=1k=-1, a rational homology disk smoothing actually exists for a singularity in the corresponding topological type by [FOW13, Theorem 5(a)], when the cross ratio of the intersection points of the node with the four arms is $9$. Our construction produces a $\mathbb{Q}$HD Stein filling in this case as well, but we do not know whether our filling is diffeomorphic to the Milnor fiber. With this in mind, our construction is an extension of the $\mathcal{C}^{4}$ family of [BS11].
@@
These graphs all correspond to sandwiched singularities (in combinatorial terms, this means that the graph can be augmented, by adding (1)(-1) vertices, to a plumbing graph that can be blown down to a smooth point). Deformation theory for this class of singularities can be understood thanks to a very attractive approach of de Jong–van Straten [dv98]: the surface singularity $(X,0)$ is encoded via the germ of a singular plane curve $(C,0)$, and all smoothings of $(X,0)$ correspond to certain picture deformations of $(C,0)$, decorated with marked points. Milnor fibers can be reconstructed directly from picture deformations. In particular, a Milnor fiber of a smoothing is a rational homology disk if and only if the number of irreducible components of $(C,0)$ matches the number of marked points of the corresponding picture deformation. The combinatorial features of the arrangement of curves and marked points in a picture deformation depend on the original dual resolution graph. In [BEK25], the first author was able to rule out rational homology disk smoothings for certain families of graphs through combinatorial analysis of possible curve and point arrangements. The family of Figure 1 was discovered in the process: combinatorial obstructions vanish, even though rational homology smoothings do not exist in this case by [WAH11, Theorem 8.6], since the graphs of Figure 1 are taut of type $(L_{2})-(J_{1})-(R_{1})$ in the sense of Laufer [LAU73, Section 2.2]. For symplectic fillings of the contact link of a sandwiched singularity, the second and third author developed an analog of the de Jong–van Straten theory [PS23, PS25], showing that all fillings arise in a similar way from certain immersed disk arrangements in $\mathbb{C}^{2}$ compatible with the germ $(C,0)$ associated to the singularity. The constructions of [PS25] use spinal open books and nearly Lefschetz fibrations; the symplectic input comes from [BH16, LHW18, LHW20, MRW24].
@@
From the de Jong–van Straten correspondence and non-existence of the corresponding $\mathbb{Q}$HD smoothings, it follows that the combinatorial arrangements producing our $\mathbb{Q}$HD fillings cannot be realized by analytic deformations of the corresponding curve germs. We do not know how to prove this deformation non-realizability by a direct argument for analytic deformation of plane curves; it would be interesting to investigate this phenomenon further.
@@ 2. Picture deformations, DJVS arrangements, and $4$-manifold constructions
A normal surface singularity $(X,0)$ is sandwiched if there exists an embedding of the tubular neighborhood of the exceptional set of its resolution $\tilde{X}$ into some blowup of $\mathbb{C}^{2}$. One can consider arbitrary resolutions here; for simplicity, we always work with the minimal good resolution, where the exceptional curves are smooth and intersect at transverse double points only. (Note that the minimal resolution is good for all rational singularities.) A good resolution can be encoded by the dual resolution graph, whose vertices correspond to the irreducible components of the exceptional divisor, and the edges record intersections between the components. Sandwiched singularities can be defined by a combinatorial condition on the dual resolution graph: a graph is sandwiched if it can be augmented by adding new end vertices with self-intersection (1)(-1) in a way that the resulting augmented graph can be blown down to a smooth point at 0. (The choice of such augmentation is generally not unique.) The (1)(-1) vertices correspond to a distinguished collection of (1)(-1) curves in a blowup of $\mathbb{C}^{2}$ with the embedding of $\tilde{X}$, so that the configuration of these (1)(-1) curves together with the original exceptional set can be completely blown down.
In [dv98], the sandwiched singularity is then encoded by an associated singular plane curve germ, as follows. For each distinguished (1)(-1) curve, fix a transverse complex disk $\tilde{C_{i}}$ through a generic point. Blow down the configuration of the exceptional curves of $(X,0)$ together with these (1)(-1) curves; the image $C_{i}$ of the disk $\tilde{C_{i}}$ under the blowdown is (a germ of) a curve in $\mathbb{C}^{2}$, possibly singular at the origin. We set $C=\cup C_{i}$ and decorate $C$ with a weight $w$, a collection of integers $w_{i}=w(C_{i})$. For each $C_{i}$, the number $w_{i}$ is the sum of multiplicities of the intersections of the image of $\tilde{C_{i}}$ with the exceptional curves during the blowdown process. (If $C_{i}$ is smooth, then $w_{i}$ is simply the number of blowdowns that the corresponding component goes through.) The original singularity can be reconstructed from the decorated germ $(C,w)$, where $C=\cup C_{i}$ and $w$ is collection of weights $w_{i}$: the corresponding blowups recover the embedding of the tubular neighborhood of the exceptional set of the resolution given by $G$ into some blowup of $\mathbb{C}^{2}$.
@@
Figure 2 shows the augmented graph, the blowdown process, and the resulting decorated germ for the graph $G_{0,1}$ of Figure 1. Each component $C_{i}$ of the germ is modeled on the cusp $y^{2}=x^{3}$ because the previous steps of the blowdown give a familiar embedded resolution of the cuspidal cubic. We also see that components $C_{i}$ and $C_{j}$ have intersection multiplicity $8$ when $i,j=1,\dots,6$, and $C_{0}$ and $C_{i}$ have multiplicity of intersection $7$ for each $i=1,\dots,6$. In general, for the graphs $G_{k,n}$ with $k\geq 0$ we get cusps $C_{1},...,C_{6+k}$ with decoration $7+k$ and pairwise intersection multiplicity $8+k$ for each pair, and instead of $C_{0}$ we have $n$ cusps $C^{\prime}_{1},\dots C^{\prime}_{n}$ with decoration $7+k+n$, with $C^{\prime}_{i}\cdot C^{\prime}_{j}=8+k+n$ for $1\leq i<j\leq n$ and $C^{\prime}_{j}\cdot C_{i}=7+k$ for $i=1,\dots,k+6$, $j=1,\dots,n$.
@@
Strictly speaking, our pictures encode topological rather than analytic data. If we consider the exceptional set of the resolution $\tilde{X}$ as an embedded configuration of complex curves, then, once the augmented graph is fixed, $C$ is a germ of a reducible complex curve in $\mathbb{C}^{2}$ (defined up to analytic equivalence). However, if we only fix the resolution graph and its augmentation, the output of the blowdown process is the topological type of a germ $C$, that is, the link $C\cap S^{3}$ of the plane curve singularity $(C,0)$ obtained by intersecting the curve $C$ with a small sphere centered at the origin. (This is a link in $S^{3}$ in the knot-theoretic sense, defined up to smooth isotopy.) By [CP04], the contact link $(Y,\xi)$ of the singularity $(X,0)$ depends only on the topological type; to generate symplectic fillings, we will only need to know the isotopy class of $C\cap S^{3}$.
@@
(C-3) the marked points $p_{j}$ are chosen on $C^{t}$, so that each self-intersection and each intersection between two irredicible components is marked; additionally, there may be free marked points on each $C_{i}$ away from the intersections,
@@
In [dv98], $p$ is the deformation of the corresponding multiplicity scheme, originally concentrated at 0 and reduced for $t>0$, but we will think of the marked points simply as a set.
The $\delta$-constant property implies that the deformation preserves the branches, and each $C^{t}_{i}$ is an immersed disk in a Milnor ball $B$, [TEI76]. With the appropriate choice of coordinates, we can assume that none of the tangent cones of the branches $C_{i}$ of the decorated germ are vertical and, moreover, all vertical tangencies of the projections $C^{t}_{i}\to\mathbb{C}_{x}$ are nondegenerate. This means that outside of self-intersections, each projection is a branched covering with simple branch points only; the degree of the covering is the multiplicity of the corresponding branch $C_{i}$ at 0. We fix a good representative of the deformation in a Milnor ball, thought of as the product $B=D_{x}\times D_{y}$ of two coordinate disks, with corners smoothed. All the constructions will take place inside this ball. We will also assume that $C\cap\partial B$ is contained in $\partial D_{x}\times D_{y}$, and think of the link of $C$ as a braid.
@@
($\Gamma$-2) All intersections $\Gamma_{i}\cap\Gamma_{k}$ and self-intersections of the components $\Gamma_{i}$ are positive transverse multiple points locally modeled on the intersection of complex lines.
($\Gamma$-3) All intersections and self-intersections are marked, and there can be additional free marked points.
@@
([dv98]) Every Milnor fiber of a deformation of a sandwiched singularity $(X,0)$ can be obtained by the DJVS construction from some picture deformation of a fixed decorated germ $(C,w)$ for $(X,0)$.
@@
([PS25]) (1) For every DJVS arrangement $(\Gamma,p)$ compatible with $(X,0)$, the 4-manifold $W_{(\Gamma,p)}$ carries a Stein structure and gives a Stein filling of the contact link $(Y,\xi)$ of $(X,0)$.
@@
Indeed, it is not hard to see that $H_{1}(W_{\Gamma,p})$ is generated by loops around $\tilde{\Gamma}_{i}$’s, $H_{2}(W)$ can be identified with the kernel of the incidence map $\mathcal{I}:\mathbb{Z}\langle{p_{j}}\rangle\to\mathbb{Z}\langle{\Gamma_{i}}\rangle$, and $H_{1}(W)$ is the cokernel of this map. We have $b_{1}=0$ because $W$ consists of 1- and 2-handles only, and $b_{1}(Y)=0$ for the link $Y=\partial W$ since the singularity is rational, c.f. [dv98, Theorem 5.2], [PS23, Section 6.1]. Thus, to produce a $\mathbb{Q}$HD filling, the arrangement $(\Gamma,p)$ must have the number of marked points matching the number of components $\Gamma_{i}$; this is rare, since typically an arrangement $\Gamma$ has a much larger number of marked points than components.
To illustrate the discussion, we describe the Milnor fiber of one of the smoothings of the singularities in our family. Consider the germ $C$ of Figure 2 and take its Scott deformation, constructed iteratively as follows. We blow up at the singular point at the origin, perform a small deformation of the strict transform so that its singularities become disjoint from the exceptional curve $E$, and the branches of the strict transforms are transverse to $E$ (equivalently, we can think of this as “shifting” the exceptional curve $E$ off the singularity in the strict transform), and then blow down $E$ to get a deformed curve $C^{\prime}$ in $\mathbb{C}^{2}$. As a result, the curve $C^{\prime}$ has a new transverse multiple point as well as the collection of singularities occurring on the strict transform of $C$ in the blowup of $\mathbb{C}^{2}$ at $p$. The procedure is repeated until all the singularities at the given stage are transverse multiple points. See [dv98, A’C75] for details, including the explanation why this procedure can be actually realized by a 1-parameter deformation. Figure 3 illustrates the Scott deformation of our decorated germ from Figure 2. The marked points on the picture deformation include all the intersections as well as the free points to match the weights on the components of the singular germ $C$. (Decorated germs look similar for all singularities corresponding to graphs $G_{0,n}$, with the number of intersection points increasing with $n$, although the parity of $n$ affects the order of curve “tails” on the right side of the picture.) For an arbitrary decorated germ, the Scott deformation produces the Artin smoothing under the de Jong–van Straten correspondence, with the Milnor fiber that is diffeomorphic to the minimal resolution of the surface singularity [dv98].